### v2

# improved consistent weighted sampling PRG pseudorandom generator

Abstract.
The distribution of keys to a given number of buckets is a fundamental task in distributed data processing and storage. A simple, fast, and therefore popular approach is to map the hash values of keys to buckets based on the remainder after dividing by the number of buckets. Unfortunately, these mappings are not stable when the number of buckets changes, which can lead to severe spikes in system resource utilization, such as network or database requests. Consistent hash algorithms can minimize remappings, but are either significantly slower than the modulo-based approach, require floating-point arithmetic, or are based on a family of hash functions rarely available in standard libraries. This paper introduces JumpBackHash, which uses only integer arithmetic and a standard pseudorandom generator. Due to its speed and simple implementation, it can safely replace the modulo-based approach to improve assignment and system stability. A production-ready Java implementation of JumpBackHash has been released as part of the Hash4j open source library.
consistent hashing, sharding, distributed storage, distributed processing
1. Introduction
Distributed data processing and storage requires a strategy how tasks or data are mapped across available resources. In more abstract terms, the strategy must be able to distribute keys to a given number of buckets. Ideally, the keys are uniformly assigned to the buckets to achieve balanced utilization of resources. If there are $n$ buckets labeled $10,1,\ldots,n-$ 1, a very popular pattern to map a given integer key $k$ to a bucket index $b$ uses the modulo operation according to
(1) $k\bmod n\rightarrow b.$
If the key $k$ is the result of some high-quality hash function (Yi, [n. d.]; Vaneev, [n. d.]; Peters, [n. d.]) and can be considered like a true random value, and the number of buckets is much smaller than the maximum possible hash value, which is usually the case for 64-bit hash values, the keys will be uniformly distributed over the buckets, such that
(2) $\Pr((k\bmod n)=b)=\frac{1}{n}.$
Unfortunately, a change in the number of buckets will also change the mapping of nearly all keys. If $n$ is increased to $n+1$, every key will be mapped to a different bucket with a probability of $\frac{n}{n+1}$ which approaches 1 for large $n$. The same is true if the number of buckets is reduced. All those reassignments would happen at the same time and cause pressure on a distributed system. For example, this might lead to timely concentrated higher network load as data needs to be shifted to other physical instances. Or, this could also result in a short-term increase of database requests to load key-specific settings or missing cache entries, which in turn may cause back pressure on data streams.
Therefore, it is critical to minimize the number of reassignments while preserving a uniform distribution. To maintain uniformity after increasing the number of buckets from $n$ to $n+1$, some keys must be shifted from the already existing $n$ buckets to the newly added $(n+1)$-th bucket with index $n$. Since the expected proportion of keys in the new bucket should be $\frac{1}{n+1}$ at the end, it is obvious that on average at least $\frac{1}{n+1}$ of the keys must switch buckets. Consistent hash algorithms achieve this lower bound and thus minimize the number of reassignments while mapping keys uniformly to a given number of buckets. For them, the probability that a key gets reassigned would be close to 0 for large $n$, because $\frac{1}{n+1}\rightarrow 0$ as $n\rightarrow\infty$. If there was no performance difference, it would be generally wise to replace pattern (1) by
$\operatorname*{ch}(k,n)\rightarrow b$
where $\operatorname*{ch}(k,n)$ denotes a consistent hash function.
Figure 1 illustrates the difference between the modulo-based approach and consistent hashing in terms of the number of reassignments when the number of buckets changes from 3 to 4. This work presents JumpBackHash, which is a simple consistent hash function with a performance comparable to the modulo-based approach (1). Table 1 provides an overview of the used notations.
Figure 1. When changing the number of buckets, some keys (letters) must be mapped to different buckets to restore balance. In contrast to the modulo operation, which puts almost all keys into different buckets, consistent hashing minimizes the expected number of reassignments (arrows).
Table 1. Notations.
Symbol Description
[…]
$\langle\ldots\rangle_{2}$ binary representation, e.g. $\langle 110\rangle_{2}=6$
$\oplus$ bitwise XOR, e.g. $\langle 1010\rangle_{2}\oplus\langle 1100\rangle_{2}=\langle 0110\rangle_{2}$
$\operatorname*{bitcount}$ number of set bits, e.g. $\operatorname*{bitcount}(\langle 10110\rangle_{2})=3$
$k$ key, typically a 64-bit hash value
$n$ number of buckets, $n>0$
$N$ some constant upper bound such that $n\leq N$ holds, see Section 2.3
$M$ some constant such that $n\leq 2^{M}$ holds, see Section 2.4
$\operatorname*{ch}(k,n)$ consistent hash function, see Section 1.1
$\mathcal{A}$ set of active indices, compare Section 2.1
$\operatorname*{Uniform}(\mathcal{S})$ uniform distribution over the set $\mathcal{S}$
$\operatorname*{Exp}(\lambda)$ exponential distribution with rate parameter $\lambda$
$\operatorname*{Geom}(p)$ geometric distribution with success probability $p$ and positive support
$\operatorname*{Bernoulli}(p)$ Bernoulli distribution with probability $p$
$\sim$ is distributed as, e.g. $\operatorname*{Bernoulli}(0.5)\sim\operatorname*{Uniform}(\{0,1\})$
$\Pr$ probability, e.g. $\Pr(\operatorname*{Bernoulli}(0.3)=1)=0.3$
$\operatorname*{\mathbb{E}}$ mean, e.g. $\operatorname*{\mathbb{E}}(\operatorname*{Uniform}(\{0,1\}))=0.5$
$\operatorname*{Var}$ variance
$\mathcal{R}$ pseudorandom generator (PRG)
$\mathcal{R}[D]$ sample from distribution $D$ using $\mathcal{R}$
$\llbracket n\rrbracket$ $\llbracket n\rrbracket:=\{0,1,2,\ldots,n-1\}$
$\mathcal{I}_{m}$ $\mathcal{I}_{m}:=\{2^{m},2^{m}+1,\ldots,2^{m+1}-1\}=\llbracket 2^{m+1}\rrbracket\setminus\llbracket 2^{m}\rrbracket$, $m\geq 0$
$\mathcal{O}$ big O notation, e.g. $\sum_{i=1}^{n}\frac{1}{i}=\mathcal{O}(\ln n)$
$\alpha_{n}$ $\alpha_{n}:=\frac{1}{n}2^{\lfloor\log_{2}(n-1)\rfloor+1}\in[1,2)$, $n\geq 2$
1.1. Consistent Hash Functions
A consistent hash function, is a deterministic function that maps a key $k$ to a bucket index from $\llbracket n\rrbracket:=\{0,1,2,\ldots,n-1\}$, where $n>0$ is the number of buckets, and that has the following two properties:
Uniformity::
The probability that a random key is assigned to a bucket is the same for all $n$ buckets which can be written as (Ioffe, 2010; Leu, 2023; Manasse et al., 2010)
$\Pr(\operatorname*{ch}(k,n)=b)=\frac{1}{n}.$
Uniform hash functions are sometimes also called balanced (Dong et al., 2023; Mendelson et al., 2021; Coluzzi et al., 2023a, b) or regular (Masson and Lee, 2024).
Monotonicity::
If the number of buckets is increased and a key is mapped to a different bucket, the bucket is always the last newly added one (Coluzzi et al., 2023b, a; Dong et al., 2023), which can be expressed as
(3) $\operatorname*{ch}(k,n)\neq\operatorname*{ch}(k,n+1)\Rightarrow\operatorname*{ch}(k,n+1)=n.$
This property is sometimes also referred to as minimal disruption (Mendelson et al., 2021) or consistency (Ioffe, 2010). In contrast, in this paper, we use the term consistency to describe the combination of monotonicity and uniformity.
While the modulo-based approach is uniform according to (2), it is not monotonic. A minimal example violating (3) is given by $k=4$ and $n=2$ when $\operatorname*{ch}(k,n):=k\bmod n$. The implication in (3) is actually an equivalence, because $\operatorname*{ch}(k,n+1)=n>\operatorname*{ch}(k,n)\in\llbracket n\rrbracket$. Therefore, the probability of reassignments for consistent hash functions is given by $\Pr(\operatorname*{ch}(k,n)\neq\operatorname*{ch}(k,n+1))=\Pr(\operatorname*{ch}(k,n+1)=n)=\frac{1}{n+1}$. Hence, uniformity and monotonicity are sufficient to minimize the expected number of reassignments.
1.2. Related Work
Interestingly, consistent hashing can be seen as a special case of consistent weighted sampling (Manasse et al., 2010), where the weight vector has a single dimension with an integer weight equal to the number of buckets $n$. The adaptation of the improved consistent weighted sampling (ICWS) algorithm (Ioffe, 2010) to our problem results in Algorithm 1 which obviously has a constant time complexity. Unfortunately, the algorithm involves expensive floating-point operations including an exponentiation and a logarithm evaluation, which makes it rather slow, especially on embedded or low-power devices without floating-point units. Nevertheless, the algorithm can be easily implemented, as it relies on an ordinary pseudorandom generator (PRG) that is seeded with the given key $k$. Standard libraries often already come with modern and reliable PRG implementations, such as SplitMix (Steele et al., 2014) in Java, that pass a series of statistical tests (Lemire, [n. d.]). These libraries sometimes even offer functions for the efficient generation of exponentially distributed random values, as required by ICWS. Although ICWS has been around for some time, it was only recently discovered as a consistent hash function when it was added to version 0.14.0 of our open source Hash4j library (Has, [n. d.]).
JumpHash (Lamping and Veach, 2014), on the other hand, became much more popular, although it was invented a few years afterwards. It is shown in Algorithm 2 and will be discussed in more detail later. As ICWS, it uses a standard PRG and floating-point arithmetic, but avoids expensive transcendental functions. Therefore, it is usually faster than ICWS, except for large $n$ due to its logarithmic time complexity. JumpHash is widely used in the meantime. For example, it is offered as hash function in the Guava Java library (Gua, [n. d.]) or in ClickHouse (Cli, [n. d.]). It is also used by Booking.com for their customer review system (Hiltpolt, 2022), by Grafana for trace and metric data distribution (Agarwal, 2020; Boucault, 2020), by VMware Greenplum since version 6 to map data values to segments (Gre, [n. d.]), by RabbitMQ since version 3.7.8 in its Consistent Hashing Exchange Plugin for message distribution (Rab, [n. d.]), or for distributed processing of vehicle license plates (Wang et al., 2023).
Recently, PowerHash was introduced as the first consistent hash algorithm with constant runtime able to beat JumpHash over the whole range of bucket numbers (Leu, 2023). Unlike JumpHash and ICWS, it relies on a family of hash functions that are less common in standard libraries and often have to be self-implemented. Since families of hash functions are used less frequently in practice, their statistical properties such as independence and uniformity are usually less tested and therefore less trustworthy than those of PRGs. A family of hash functions is also used by the even more recent FlipHash algorithm (Masson and Lee, 2024), which is similar to PowerHash but avoids floating-point operations.
function ICWS($k$, $n$)
       $\mathcal{R}\leftarrow$ initialize PRG with seed $k$
       $U\leftarrow\mathcal{R}[\operatorname*{Uniform}([0,1))]$
       $G\leftarrow\mathcal{R}[\operatorname*{Exp}(1)]+\mathcal{R}[\operatorname*{Exp}(1)]$
        $\triangleright$ $G$ follows a $\operatorname*{Gamma}(2,1)$ distribution
       return $\min(\lfloor\exp(G\cdot(\lfloor\ln(n)/G+U\rfloor-U))\rfloor,n-1)$
        $\triangleright$ truncate at $n-1$ to protect against numerical errors
Algorithm 1 Specialization of improved consistent weighted sampling (ICWS) (Ioffe, 2010) to map a key $k$ consistently to $\llbracket n\rrbracket:=\{0,1,2,\ldots,n-1\}$.
function JumpHash($k$, $n$)
       $\mathcal{R}\leftarrow$ initialize PRG with seed $k$
       $b\leftarrow-1$
       $b^{\prime}\leftarrow 0$
       while $b^{\prime}<n$ do
             $b\leftarrow b^{\prime}$
             $U\leftarrow\mathcal{R}[\operatorname*{Uniform}([0,1))]$
             $b^{\prime}\leftarrow\lfloor(b+1)/U\rfloor$
              $\triangleright$ determine next larger active index, compare (7)
      return $b$
Algorithm 2 The JumpHash algorithm (Lamping and Veach, 2014) consistently maps $k$ to the integer range $\llbracket n\rrbracket:=\{0,1,2,\ldots,n-1\}$.
Consistent hashing as considered in this work, allows removing buckets only in the reverse order as they were added. A lot of research was done on the more general problem, where buckets can be arbitrarily added and removed (Thaler and Ravishankar, 1998; Appleton and O’Reilly, 2015; Karger et al., 1997; Dong et al., 2023; Eisenbud et al., 2016; Nakatani, 2021; Mendelson et al., 2021; Mirrokni et al., 2018; Chen et al., 2021; Coluzzi et al., 2023a; Zeng et al., 2022). However, they are typically significantly slower, require some state, or are not optimal with respect to the number of reassignments. A recently presented algorithm, MementoHash (Coluzzi et al., 2023b) builds upon JumpHash and therefore demonstrates how any consistent hash algorithm can be extended to support removals in arbitrary order.
1.3. Our Contributions
We introduce JumpBackHash, a fast consistent hash algorithm with expected constant runtime, which also works completely without floating-point operations. In contrast to PowerHash (Leu, 2023) or FlipHash (Masson and Lee, 2024) which require a family of hash functions, JumpBackHash, like JumpHash (Lamping and Veach, 2014), only uses a standard pseudorandom generator (PRG).
To derive JumpBackHash, we use the concept of active indices, which has already been used for consistent weighted sampling (Manasse et al., 2010; Ioffe, 2010). This new perspective allows a better understanding of JumpHash and leads us after some stepwise optimizations to JumpBackHash. We also perform a theoretical runtime analysis, which is empirically verified together with uniformity and monotonicity.
2. Methodology
A simple way to define a consistent hash function is
(4) $\operatorname*{ch}(k,n):=\operatorname*{arg\,min}_{b\in\llbracket n\rrbracket}R_{b},$
where $R_{0},R_{1},R_{2},\ldots$ is a sequence of independent and identically distributed values obtained by a PRG with seed $k$ and a support large enough to ignore accidental collisions in practice. This function is uniform as the random values $R_{0},R_{1},\ldots,R_{n-1}$ have the same chance of being the minimum. It is also monotonic, because if the number of buckets is increased from $n$ to $n+1$, the mapping remains stable, except for the case where $R_{n}$ is the new minimum and thus the result changes to $n$.
Unfortunately, this approach is not very efficient as it takes $\mathcal{O}(n)$ time to generate all relevant random values $R_{b}$ with $b\in\llbracket n\rrbracket$. The order in which the random values $R_{b}$ are drawn using the PRG is not important as long as it is well defined and does not depend on $n$. For example, if we know an upper bound $N$ such that $n\leq N$ is always satisfied, we could draw the random values in reverse order according to $R_{N-1},R_{N-2},\ldots,R_{1},R_{0}$. Even though this would lead to a much worse time complexity of $\mathcal{O}(N)$ if $n\ll N$, (4) would still be a valid consistent hash function.
2.1. Active Indices
To construct a more efficient consistent hash algorithm based on (4), we use the concept of active indices (Manasse et al., 2010; Ioffe, 2010). A bucket index $b$ is an active index if it satisfies
(5) $R_{b}=\min_{b^{\prime}\in\llbracket b+1\rrbracket}R_{b^{\prime}}.$
By definition, bucket index 0 is always an active index. For $b>0$ this definition is equivalent to $R_{b}<\min_{b^{\prime}\in\llbracket b\rrbracket}R_{b^{\prime}}$. According to (4) only active indices are candidates for the mapping. Let $\mathcal{A}$ be the set of active indices, then (4) can be simply expressed as
(6) $\operatorname*{ch}(k,n):=\max(\mathcal{A}\cap\llbracket n\rrbracket),$
If we find an equivalent statistical process to generate the active indices in order and directly, without the detour of generating the random values $R_{b}$ for each individual bucket index, we could save a lot of computation time.
2.2. JumpHash
Generating active indices in ascending order is exactly what JumpHash does, even though the algorithm was derived and analyzed without the concept of active indices (Lamping and Veach, 2014). Given an active index $A$, the probability that $A+1$ is the next larger active index equals the probability that $R_{A+1}$ is the smallest among the first $A+2$ random values $\{R_{b}\}_{b\in\llbracket A+2\rrbracket}$ which occurs with a probability of $\frac{1}{A+2}$. The probability that $A+2$ is the next active index relative to $A$ is given by the probability that $A+1$ is not an active index multiplied by the probability that $R_{A+2}$ is the smallest value, which is therefore $(1-\frac{1}{A+2})\frac{1}{A+3}=\frac{A+1}{(A+2)(A+3)}$. In general, the probability that $A+\Delta$ is the next active index relative to $A$ is given by $(\prod_{i=2}^{\Delta}1-\frac{1}{A+i})\frac{1}{A+\Delta+1}=\frac{A+1}{(A+\Delta)(A+\Delta+1)}$.
Hence, the active indices can be generated directly and in ascending order using the recursion
(7) $A_{l+1}=\lfloor(A_{l}+1)/U\rfloor\qquad\text{with $U\sim\operatorname*{Uniform}([0,1))$ and $A_{0}=0$,}$
because $A_{l+1}=A_{l}+\Delta$ holds if $U\in(\frac{A_{l}+1}{A_{l}+\Delta+1},\frac{A_{l}+1}{A_{l}+\Delta}]$, which occurs with a probability of $\frac{A_{l}+1}{A_{l}+\Delta}-\frac{A_{l}+1}{A_{l}+\Delta+1}=\frac{A_{l}+1}{(A_{l}+\Delta)(A_{l}+\Delta+1)}$ as required. The recursion together with (6) leads directly to the JumpHash algorithm as shown in Algorithm 2.
The runtime is given by the number of consumed random values $T_{\text{JH}}$ which also corresponds to the number of active indices that need to be generated until the first active index greater than or equal to $n$ is found. As the very first active index is always 0 and does not have to be generated, $T_{\text{JH}}$ is also the same as the number of active indices in $\llbracket n\rrbracket$. The probability that bucket index $b$ is active is $\frac{1}{b+1}$, as this is the probability that the associated random value $R_{b}$ is the smallest among the first $b+1$ random values $\{R_{b^{\prime}}\}_{b^{\prime}\in\llbracket b+1\rrbracket}$. Therefore, $T_{\text{JH}}$ is distributed as
$T_{\text{JH}}\sim\sum_{b=0}^{n-1}\operatorname*{Bernoulli}\!\left(\frac{1}{b+1}\right)\sim\sum_{i=1}^{n}\operatorname*{Bernoulli}\!\left(\frac{1}{i}\right),$
where $\operatorname*{Bernoulli}(p)$ denotes a Bernoulli random variable with probability $p$. When using
$\operatorname*{\mathbb{E}}\!\left(\sum_{i=1}^{n}K_{i}\right)$ $=\sum_{i=1}^{n}\operatorname*{\mathbb{E}}(K_{i}),$
$\operatorname*{Var}\!\left(\sum_{i=1}^{n}K_{i}\right)$ $=\sum_{i=1}^{n}\operatorname*{Var}(K_{i})$
for independent $K_{i}$ together with the known expressions for Bernoulli distributed random variables
(8) $\operatorname*{\mathbb{E}}(\operatorname*{Bernoulli}(p))$ $=p,$
(9) $\operatorname*{Var}(\operatorname*{Bernoulli}(p))$ $=p(1-p),$
the mean and the variance of $T_{\text{JBH}}$ can be derived as
(10) $\operatorname*{\mathbb{E}}$ $(T_{\text{JH}})$ $=\sum_{i=1}^{n}\frac{1}{i}$ $\leq 1+\ln n,$
(11) $\operatorname*{Var}$ $(T_{\text{JH}})$ $=\sum_{i=1}^{n}\frac{1}{i}\left(1-\frac{1}{i}\right)=\left(\sum_{i=1}^{n}\frac{1}{i}\right)-\left(\sum_{i=1}^{n}\frac{1}{i^{2}}\right)$ $\leq\ln n.$
This also shows the logarithmic runtime complexity $\mathcal{O}(\ln n)$ of JumpHash.
2.3. Jumping Backwards
The way JumpHash generates active indices (7) is not unique. If an upper bound $N$ is known for the number of buckets $n\leq N$, which is always the case in practice as standard integer types are limited, it is also possible to generate active indices in descending order using the recursion
(12) $A_{l+1}\sim\operatorname*{Uniform}(\llbracket A_{l}\rrbracket)\qquad\text{with$l\geq 0$ and $A_{0}=N$.}$
The rationale behind this approach is very simple. The next smallest active index is given by the smallest random value from $\{R_{b}\}_{b\in\llbracket A_{l}\rrbracket}$. As, by definition, the values in this set are all identically distributed, they all have the same chance of being the smallest. As a result, the next smaller active index is uniformly distributed over $\llbracket A_{l}\rrbracket$. Based on this recursion (12) another consistent hash algorithm can be constructed as shown by Algorithm 3.
function JumpingBackwards($k$, $n$)
       $\mathcal{R}\leftarrow$ initialize random generator with seed $k$
       $b\leftarrow N$
       repeat
             $b\leftarrow\mathcal{R}[\operatorname*{Uniform}(\llbracket b\rrbracket)]$
              $\triangleright$ determine next smaller active index, compare (12)
      until $b<n$
      return $b$
Algorithm 3 Consistent hashing by jumping backwards, $n\leq N$.
In contrast to JumpHash, this approach does not involve floating-point operations. However, it requires sampling from an integer range (O’Neil, 2018; Lemire, 2019). Another downside is that all active indices in $[n,N)$ must be generated. If $n\ll N$, which is usually the case, the runtime complexity would be $\mathcal{O}(\log N)$, which is worse than the $\mathcal{O}(\log n)$ of JumpHash. While we prefer the absence of floating-point operations, we need to make the generation of active indices in descending order much more efficient.
2.4. Improved Active Index Generation
To construct a more efficient process to generate active indices in descending order we assume without loss of generality that the maximum number of buckets is a power of two $N=2^{M}$ and divide the set of bucket indices $\llbracket N\rrbracket$ into disjoint intervals $\mathcal{I}_{m}:=\{2^{m},2^{m}+1,\ldots,2^{m+1}-1\}$ according to
$\llbracket N\rrbracket=\{0\}\cup\bigcup_{m=0}^{M-1}\mathcal{I}_{m}.$
A good reason to consider intervals with power-of-two cardinalities is that uniform integer values can be efficiently sampled from $\mathcal{I}_{m}$ using exactly $m$ random bits. The intervals $\mathcal{I}_{m}$ can be ordered as all values of $\mathcal{I}_{m}$ are smaller than that of $\mathcal{I}_{m+1}$ and also the cardinality of $\mathcal{I}_{m}$ is smaller than that of $\mathcal{I}_{m+1}$. If we compare those intervals in the following, we always refer to this ordering. In particular, this allows us to say that $\mathcal{I}_{m}$ is smaller than $\mathcal{I}_{m^{\prime}}$ if $m<m^{\prime}$.
To obtain active indices in descending order, we process the intervals $\mathcal{I}_{m}$ with $m\in\llbracket M\rrbracket$ in descending order and generate all active indices therein also in descending order. The last active index is always 0 by definition (5). The probability that $\mathcal{I}_{m}$ contains an active index is 50% as the next smaller random value belongs either to $\{R_{b}\}_{b\in\llbracket 2^{m}\rrbracket}$ or $\{R_{b}\}_{b\in\mathcal{I}_{m}}$ with the same probability. This decision requires a single random bit $X_{m}\in\{0,1\}$. If $X_{m}=0$, which occurs in 50% of all cases, we know that there is no active index in $\mathcal{I}_{m}$ and we can continue processing the next smaller interval. Otherwise, if $X_{m}=1$, there is at least one active index, but we still need to find out which indices from $\mathcal{I}_{m}$ are actually active.
As all indices $b\in\mathcal{I}_{m}$ have the same chance that their associated random value $R_{b}$ is the smallest among them, the largest active index in $\mathcal{I}_{m}$ can simply be generated by drawing a uniform random value $Y_{m}$ from $\mathcal{I}_{m}$. It is possible that $\mathcal{I}_{m}$ contains further (smaller) active indices in the interval. As the next smaller active index (not necessarily in $\mathcal{I}_{m}$) is uniformly distributed over $\llbracket Y_{m}\rrbracket$, we sample a uniform random value $Z_{m,0}$ from $\llbracket Y_{m}\rrbracket$ analogous to (12). If $Z_{m,0}\notin\mathcal{I}_{m}$, $Z_{m,0}$ is rejected as active index and the process is continued with the next smaller interval $\mathcal{I}_{m-1}$. Otherwise, $Z_{m,0}$ is the next smaller active index within $\mathcal{I}_{m}$, and we iteratively generate further random values according to
(13) $Z_{m,v}\sim\operatorname*{Uniform}(\llbracket Z_{m,v-1}\rrbracket)\qquad\text{with $v\geq 1$},$
as long as they are within $\mathcal{I}_{m}$, to get all active indices of that interval.
The whole procedure to generate active indices in descending order, which is statistically equivalent to Algorithm 3, is summarized in Algorithm 4. While $X_{m}$ and $Y_{m}$ are already very convenient to sample, $Z_{m,v}$ needs to be sampled from a non-power-of-two interval $\llbracket b\rrbracket$. One possibility would be to use Lemire’s method (Lemire, 2019). A simpler and also unbiased approach that fits better in our case is to sample from the smallest enclosing power-of-two interval, which is by definition $\llbracket 2^{m+1}\rrbracket$ for $\mathcal{I}_{m}$, according to
(14) $Z_{m,v}\sim\operatorname*{Uniform}(\llbracket 2^{m+1}\rrbracket)\qquad\text{with $v\geq 1$}.$
As the sequence $(Z_{m,v})_{v\geq 0}$ is no longer strictly monotonically decreasing, we must ignore values that are not smaller than all previous values to ensure equivalent behavior, when replacing (13) in the inner loop of Algorithm 4, labeled as option 1, by (14), labeled as option 2. Fortunately, the inner loop takes care of this without any further modifications. If there is a value that fulfills one of the two stop conditions, but there was already another smaller or equal value before, this other value would have stopped the loop earlier. We continue with option 2, as all required random values are then either Boolean values or values from a power-of-two interval and can therefore be generated easily and efficiently by taking a corresponding number of random bits.
function JumpingBackwardsImproved($k$, $n$)
       $\mathcal{R}\leftarrow$ initialize PRG with seed $k$
       for $m\leftarrow M-1$ to 00 do $\triangleright$ process intervals $\mathcal{I}_{M-1},\mathcal{I}_{M-2},\ldots,\mathcal{I}_{0}$ in descending order
             $X_{m}\leftarrow\mathcal{R}[\operatorname*{Uniform}(\{0,1\})]$
              $\triangleright$ generate a random bit
             if $X_{m}=0$ then continue
              $\triangleright$ continue with next interval if $\mathcal{I}_{m}$ does not contain any active indices $Y_{m}\leftarrow 2^{m}+\mathcal{R}[\operatorname*{Uniform}(\llbracket 2^{m}\rrbracket)]$
              $\triangleright$ sample largest active index $Y_{m}$ from $\mathcal{I}_{m}$
             $b\leftarrow Y_{m}$
             for $v\leftarrow 0$ to $\infty$ do
                   if $b<n$ then return $b$
                    $\triangleright$ the first active index in $\mathcal{I}_{m}$ smaller than $n$ is returned $Z_{m,v}\leftarrow\mathcal{R}[\operatorname*{Uniform}(\llbracket b\rrbracket)]$
                    $\triangleright$ option 1, compare (13)
                   $Z_{m,v}\leftarrow\mathcal{R}[\operatorname*{Uniform}(\llbracket 2^{m+1}\rrbracket)]$
                    $\triangleright$ option 2, compare (14)
                   $b\leftarrow Z_{m,v}$
                   if $b<2^{m}$ then break
                    $\triangleright$ stop if there are no further active indices in $\mathcal{I}_{m}$
      return 00
Algorithm 4 Consistent hashing based on a more efficient generation of active indices in reverse order, $n\leq 2^{M}$.
2.5. Saving Random Values
Algorithm 4 requires many random values. For every interval $\mathcal{I}_{m}$ a bit $X_{m}\in\{0,1\}$, an integer $Y_{m}\in\mathcal{I}_{m}$, and an integer sequence $(Z_{m,v})_{v\geq 0}$ with $Z_{m,v}\in\llbracket 2^{m+1}\rrbracket$ need to be sampled when using option 2. The sequences $(Z_{m,v})_{v\geq 0}$ are finite, as it is sufficient to generate values up to the first value that is smaller than $2^{m}$ and that causes leaving the inner loop of Algorithm 4 in any case. We are free to change the order in which those independent random values and sequences are generated using the same PRG without changing the statistical properties of Algorithm 4. However, the order must be well-defined and not depend on $n$ in order to preserve monotonicity.
As first optimization, we generate all $M$ random bits $X_{m}$ at once by drawing a random $M$-bit integer $U=\langle X_{M-1}X_{M-2}\ldots X_{0}\rangle_{2}\in\llbracket 2^{M}\rrbracket$. Knowing all those bits at once allows quick identification of intervals that may contain active indices that are relevant for a given $n$. The indices of all set bits of $U$ can be computed in descending order according to the recursion
(15) $m_{i+1}=\lfloor\log_{2}(U\bmod 2^{m_{i}})\rfloor$
as long as the argument of the binary logarithm is nonzero. If it is zero there are no further intervals $\mathcal{I}_{m}$ with active indices. As only intervals $\mathcal{I}_{m}$ are relevant that have the corresponding bit $X_{m}$ set and that have an overlap with $\llbracket n\rrbracket$, we can start our recursion with interval index
(16) $m_{0}=\lfloor\log_{2}(n-1)\rfloor+1\qquad\text{for $n\geq 2$}$
which is the solution of the inequality
$2^{m_{0}-1}<n\leq 2^{m_{0}}.$
Thus, $\mathcal{I}_{m_{0}}$ is the smallest interval without overlap with $\llbracket n\rrbracket$.
Among the set of intervals $\{\mathcal{I}_{m}\}_{m\in\llbracket M\rrbracket}$, at most two need to be processed as $\mathcal{I}_{{m_{1}}}$ and $\mathcal{I}_{{m_{2}}}$ (if they exist) are the largest two intervals that contain at least one active index and overlap with $\llbracket n\rrbracket$. As $\mathcal{I}_{{m_{2}}}\subseteq\llbracket n\rrbracket$ by definition, $\mathcal{I}_{{m_{2}}}$ must contain an active index that is also smaller than $n$. Therefore, the active indices of smaller intervals $\mathcal{I}_{m}$ with $m<{m_{2}}$ are not relevant. Furthermore, within $\mathcal{I}_{{m_{2}}}$ just the largest active index is needed, which means that the random sequence $(Z_{{m_{2}},v})_{v\geq 0}$ is never used. The only sequence that may be required for generating the largest active index smaller than $n$ is $(Z_{{m_{1}},v})_{v\geq 0}$ as the active indices of $\mathcal{I}_{m_{1}}$ are not necessarily smaller than $n$.
Since no more than one of the sequences $(Z_{0,v})_{v\geq 0}$, $(Z_{1,v})_{v\geq 0},\ldots,(Z_{M-1,v})_{v\geq 0}$ is used at the same time, it is not necessary that those sequences are statistically independent to ensure uniformity. Therefore, we are allowed to generate a single sequence $(W_{v})_{v\geq 0}$ with $W_{v}\in\llbracket 2^{M}\rrbracket$ instead and set
(17) $Z_{m,v}=W_{v}\bmod 2^{m+1}.$
It should be noted that with this modification, the generation of active indices is no longer statistically equivalent to (4). However, uniformity is preserved as the consumed random values are still independent. Monotonicity is also still ensured, as the active indices are still generated in descending order.
We can also reduce the randomness needed for the values $Y_{m}$. As only active indices of at most the two intervals $\mathcal{I}_{m_{1}}$ and $\mathcal{I}_{m_{2}}$ are relevant, only the two random integers $Y_{m_{1}}$ and $Y_{m_{2}}$ are needed at the same time. Thus, using the fact that $1+\sum_{i=0}^{m_{2}}X_{i}=\sum_{i=0}^{m_{1}}X_{i}$, we can set
(18) $Y_{m}=2^{m}+(V_{s}\bmod 2^{m})\qquad\text{with $\textstyle s=(\sum_{i=0}^{m}X_{i})\bmod 2$}$
with $V_{0},V_{1}\in\llbracket 2^{M}\rrbracket$ being two independent random values, as this is sufficient for the independence of $Y_{m_{1}}$ and $Y_{m_{2}}$.
If we look at the random variables $V_{0}$ and $V_{1}$, the ${m_{1}}$ least significant bits are used from one and the least significant ${m_{2}}$ bits from the other. As a consequence, the most significant $M-{m_{2}}$ bits of either $V_{0}$ or $V_{1}$ are still unused. After combining both using the XOR-operation, the most significant $M-{m_{2}}$ bits of $V_{0}\oplus V_{1}$ are therefore statistically independent of $Y_{m_{1}}$ and $Y_{m_{2}}$. This allows us to use $V_{0}$ and $V_{1}$ also for $U$ by setting
(19) $U=V_{0}\oplus V_{1},$
because we only need the random bits $X_{M-1},\ldots,X_{{m_{2}}}$ to determine ${m_{1}}$ and ${m_{2}}$ and therefore never consume more than the leading $M-{m_{2}}$ bits of $U$.
In summary, it is sufficient to generate $V_{0}$ and $V_{1}$ followed by the sequence $(W_{v})_{v\geq 0}$ to generate all needed random values without breaking uniformity and monotonicity. Therefore, all these values can be taken from a single random sequence which allows to use a standard PRG.
2.6. JumpBackHash
function JumpBackHash($k$, $n$)
       if $n\leq 1$ then return 0
       $\mathcal{R}\leftarrow$ initialize PRG with seed $k$
       $V_{0}\leftarrow\mathcal{R}[\operatorname*{Uniform}(\llbracket 2^{M}\rrbracket)]$
       $V_{1}\leftarrow\mathcal{R}[\operatorname*{Uniform}(\llbracket 2^{M}\rrbracket)]$
       $U\leftarrow V_{0}\oplus V_{1}$
        $\triangleright$ compare (19)
       $m\leftarrow\lfloor\log_{2}(n-1)\rfloor+1$
        $\triangleright$ equals $m_{0}$, see (16)
       $u\leftarrow U\bmod 2^{m}$
       while $u\neq 0$ do $\triangleright$ iterate over intervals $\mathcal{I}_{m_{1}},\mathcal{I}_{m_{2}},\ldots$ that contain at least one active index
             $m\leftarrow\lfloor\log_{2}u\rfloor$
              $\triangleright$ equals $m_{i}$ in $i$-th iteration, compare (15)
             $s\leftarrow\operatorname*{bitcount}(u)\bmod 2$
              $\triangleright$ determine parity of set bits of $u$
             $b\leftarrow 2^{m}+(V_{s}\bmod 2^{m})$
              $\triangleright$ corresponds to $Y_{m}$, compare (18)
             for $v\leftarrow 0$ to $\infty$ do $\triangleright$ $v$ could be eliminated but is kept for better comparability with Algorithm 4
                   if $b<n$ then return $b$
                    $\triangleright$ condition is always satisfied latest in the second while-iteration $W_{v}\leftarrow\mathcal{R}[\operatorname*{Uniform}(\llbracket 2^{M}\rrbracket)]$
                   $b\leftarrow W_{v}\bmod 2^{m+1}$
                    $\triangleright$ corresponds to $Z_{m,v}$, compare (17)
                   if $b<2^{m}$ then break
                    $\triangleright$ stop if there are no further active indices in $\mathcal{I}_{m}$
            $u\leftarrow u\oplus 2^{m}$
              $\triangleright$ has the same effect as $u\leftarrow u\bmod 2^{m}$
      return 00
Algorithm 5 JumpBackHash, $n\leq 2^{M}$.
Applying all these optimizations to Algorithm 4 with option 2 finally leads to JumpBackHash as shown in Algorithm 5. The variables $U$ and $W_{v}$ as well as the loop variable $v$ could be eliminated, but are kept for better comprehensibility and better comparability with Algorithm 4. The while loop iterates over the intervals $\mathcal{I}_{m_{1}},\mathcal{I}_{m_{2}},\ldots$ according to (15). Since we know that $\mathcal{I}_{m_{2}}$, if it exists, must contain an active index smaller than $n$, the condition for the return statement within the inner loop for $\mathcal{I}_{m_{2}}$ is always fulfilled. This means that there will never be more than two iterations of the while loop.
Algorithm 6 shows how the algorithm can be implemented with a few lines of code in Java. Expressions of kind $\lfloor\log_{2}u\rfloor$ can be replaced by 31-numberOfLeadingZeros(u) if $u$ is a positive 32-bit integer. Furthermore, bit-shift operations are well-defined in Java. For example, shifting by $-29$ bits is the same as shifting by 3 bits. Therefore, shifting by 31-numberOfLeadingZeros(u) is the same as shifting by -1-numberOfLeadingZeros(u) or by ~numberOfLeadingZeros(u) when using Java’s bitwise complement operator. The numberOfLeadingZeros and bitCount Java functions are intrinsic candidates. They are usually translated directly into specific CPU instructions and are therefore very fast.
As PRGs typically produce 64-bit values, we can use them to get two 32-bit random integers at once. Therefore, Algorithm 6 uses the least and most significant 32 bits of the first 64-bit random value as $V_{0}$ and $V_{1}$, respectively. Also the inner while loop uses a 64-bit random value to always generate two elements of the sequence $(W_{v})_{v\geq 0}$ at once. If two random values are always generated at once, we refer to JumpBackHash* in the following.
int jumpBackHash(long k, int n) { if (n <= 1) return 0; randomGenerator.resetWithSeed(k); long v = randomGenerator.nextLong(); int u = (int) (v ^ (v >>> 32)) & (~0 >>> numberOfLeadingZeros(n - 1)); while (u != 0) { int q = 1 << ~numberOfLeadingZeros(u); // q = 2^m int b = q + ((int) (v >>> (bitCount(u) << 5)) & (q - 1)); while (true) { if (b < n) return b; long w = randomGenerator.nextLong(); b = (int) w & ((q << 1) - 1); if (b < q) break; if (b < n) return b; b = (int) (w >>> 32) & ((q << 1) - 1); if (b < q) break; } u ^= q; } return 0; }
Algorithm 6 Java implementation of JumpBackHash using a 64-bit pseudorandom generator.
2.7. Runtime Analysis
The runtime of Algorithm 5 scales with the number of consumed random values. Therefore, we analyze the distribution of the number of consumed random values which corresponds to the number of PRG calls. As the case $n=1$ does not require any random values, we can focus on the case $n\geq 2$.
First, we need two random values for $V_{0}$ and $V_{1}$ (compare Section 2.5). Further random values are only needed if the largest active index in $\mathcal{I}_{m_{1}}$ is greater than or equal to $n$. This can only happen, if $n\in\mathcal{I}_{m_{1}}$, and thus if $m_{1}=m_{0}-1$. The probability that $\mathcal{I}_{m_{0}-1}$ contains an active index at all is $\frac{1}{2}$. As the largest active index would be uniformly distributed over $\mathcal{I}_{m_{0}-1}$, the probability that $\mathcal{I}_{m_{0}-1}$ contains an active index that is larger than or equal to $n$, or equivalently from $[n,2^{m_{0}})$, is given by $\frac{1}{2}\cdot\frac{2^{m_{0}}-n}{2^{m_{0}-1}}=1-\alpha_{n}^{-1}$ with
(20) $\alpha_{n}:=\frac{2^{m_{0}}}{n}=\frac{2^{\lfloor\log_{2}(n-1)\rfloor+1}}{n}\in[1,2).$
Only if this is the case we need to start generating the random sequence $(W_{v})_{v\geq 0}$ to find the largest active index within $\mathcal{I}_{m_{0}-1}$ that is actually smaller than $n$. The inner loop is left, either via the break or the return statement, as soon as $W_{v}\bmod 2^{m_{1}+1}=W_{v}\bmod 2^{m_{0}}<n$ which occurs with probability $\frac{n}{2^{m_{0}}}=\alpha_{n}^{-1}$. Therefore, the number of random values consumed within the inner loop follows a geometric distribution with positive support and a success probability of $\alpha_{n}^{-1}$.
In summary, the number of random values $T_{\text{JBH}}$ that need to be generated is distributed like
(21) $T_{\text{JBH}}\sim 2+\operatorname*{Bernoulli}(1-\alpha_{n}^{-1})\cdot\operatorname*{Geom}(\alpha_{n}^{-1})\qquad\text{for $n\geq 2$}.$
Using the identities for the mean and variance of the product of independent $K$ and $L$ (Goodman, 1960)
$\operatorname*{\mathbb{E}}(KL)$ $=\operatorname*{\mathbb{E}}(K)\operatorname*{\mathbb{E}}(L),$
$\operatorname*{Var}(KL)$ $=(\operatorname*{\mathbb{E}}(K))^{2}\operatorname*{Var}(L)+(\operatorname*{\mathbb{E}}(L))^{2}\operatorname*{Var}(K)+\operatorname*{Var}(K)\operatorname*{Var}(L)$
together with (8), (9), and the known expressions for geometrically distributed random variables
$\operatorname*{\mathbb{E}}(\operatorname*{Geom}(p))$ $=\frac{1}{p},$
$\operatorname*{Var}(\operatorname*{Geom}(p))$ $=\frac{1-p}{p^{2}},$
the mean and the variance of $T_{\text{JBH}}$ can be derived as
(22) $\operatorname*{\mathbb{E}}$ $(T_{\text{JBH}})$ $=1+\alpha_{n}$ $\in[2,3),$
(23) $\operatorname*{Var}$ $(T_{\text{JBH}})$ $=(\alpha_{n}-1)\alpha_{n}$ $\in[0,2).$
Hence, on average, JumpBackHash requires less than 3 random values with a variance that is always less than 2.
For JumpBackHash*, when two random values are always generated at once as in Algorithm 6, the probability that one of both random values satisfies any stop condition within the inner loop is given by $1-(1-\alpha_{n}^{-1})(1-\alpha_{n}^{-1})=\alpha_{n}^{-1}(2-\alpha_{n}^{-1})$. Therefore, as also just a single random value is needed at the beginning, the number of generated random values is distributed as
(24) $T_{\text{JBH*}}\sim 1+\operatorname*{Bernoulli}(1-\alpha_{n}^{-1})\cdot\operatorname*{Geom}(\alpha_{n}^{-1}(2-\alpha_{n}^{-1}))\qquad\text{for $n\geq 2$}$
with mean and variance given by
(25) $\operatorname*{\mathbb{E}}$ $*\displaystyle(T_{\text{JBH*}})$ $=1+\frac{(\alpha_{n}-1)\alpha_{n}}{2\alpha_{n}-1}$ $\in\textstyle[1,{\frac{5}{3}}),$
(26) $\operatorname*{Var}$ $*\displaystyle(T_{\text{JBH*}})$ $=\frac{\alpha_{n}(\alpha_{n}-1)(\alpha_{n}^{2}-\alpha_{n}+1)}{(2\alpha_{n}-1)^{2}}$ $\in\textstyle[0,{\frac{2}{3}}).$
Hence, Algorithm 6 consumes on average at most $\frac{5}{3}\approx 1.667$ 64-bit random integers.
$\alpha_{n}$ is minimized and equal to 1 if $n$ is exactly a power of two. The maxima of $\alpha_{n}$ occur whenever the number of buckets are powers of two incremented by 1, $n=2^{i}+1$ and approach 2 as $n\rightarrow\infty$. This leads to a sawtooth pattern on the log-scale. As the expressions (22), (23), (25), and (26) are all monotonically increasing functions of $\alpha_{n}$ over $[1,2)$, they replicate this sawtooth pattern. In contrast to JumpHash, where not only the mean (10) but also the variance (11) increases logarithmically, both are bounded in case of JumpBackHash.
3. Experimental Verification
To validate JumpBackHash, we have performed a series of tests. We used the SplitMix random generator (Steele et al., 2014) which produces a sequence of 64-bit random integers, is known to have good statistical properties (Lemire, [n. d.]), and is also available in the Java standard library as SplittableRandom. The source code to run all tests and to reproduce the presented results has been published at https://github.com/dynatrace-research/jumpbackhash-paper. The tested JumpBackHash implementation is equivalent to the one included in our Hash4j Java open source library since version 0.17.0 (Has, [n. d.]).
Figure 2. The measured mean and variance of the number of consumed random values for JumpHash and JumpBackHash. JumpBackHash* corresponds to the case when a single random value is split into two as in Algorithm 6.
3.1. Consistency
[…]
Testing uniformity is more sophisticated as it requires statistical means. Our strategy was to sample many bucket indices by calling the consistent hash function for many different keys. In our test we always used 1 million randomly generated keys. If uniformity holds, we expect that the bucket index distribution is uniform. For small numbers of buckets, as long as it is much smaller than the number of sampled keys, this can be verified by applying the G-test (McDonald, 2014). In this way we checked the uniformity for all $n\in\{1,2,\ldots,$1000$\}$. Additionally, we also investigated if uniformity holds for large $n$ close to $12^{31}-$ 1 which is the largest possible positive 32-bit integer value in Java. In particular, we have tested the values $n\in\{2^{31}-1,2^{31}-2,2^{30}+1,2^{30},2^{30}-1,3\cdot 2^{28},2^{29}+1,2^{29},2^{29}-1,3\cdot 2^{27},2^{28}+1,2^{28},2^{28}-1\}$ using the Kolmogorov-Smirnov test (Kanji, 2006), as the number of occurrences per possible outcome would be too low for a valid application of the G-test.
3.2. Random Value Consumption
We also checked the analytical expressions for the number of consumed random values, which is a measure of the runtime. For this purpose, we ran simulations for different $n$. To cover many orders of magnitude, we picked values for $n$ according to the recursion
(27) $n_{i+1}=\lfloor 0.999\cdot n_{i}\rfloor\qquad\text{with $n_{0}=10^{6}$}.$
This results in 7482 positive integer values that are roughly uniformly distributed on the logarithmic scale over the range $[1,10^{6}]$.
For each value of $n$ we computed the bucket index for 10 million random keys and counted how often the pseudorandom generator was called. The corresponding sample mean and variance are shown in Figure 2 for JumpHash and JumpBackHash. We also considered JumpBackHash* which splits a random value into two and which corresponds to Algorithm 6. All simulated data points for the mean did not differ from the theoretical prediction by more than 0.0036 in absolute values. Likewise, the observed difference between the empirical and the theoretical variance was always less than 0.025. We have therefore omitted plotting the theoretical curves as they look identical and overlap perfectly, which also confirms the correctness of our theoretical analysis.
3.3. Performance Measurements
Figure 3. Benchmark results on an Amazon EC2 c5.metal instance with Intel Xeon Platinum 8275CL CPU and processor P-state set to 1. Unlike Figure 2, a logarithmic scale is used for the y-axis.
We evaluated the performance of JumpBackHash on an Amazon EC2 c5.metal instance with Intel Xeon Platinum 8275CL CPU. To reduce the impact of variable processor frequencies, Turbo Boost was disabled by setting the processor P-state to 1 (Ama, [n. d.]). The benchmarks were implemented using the Java Microbenchmark Harness (JMH) toolkit and were executed using OpenJDK 21.0.2.
[…]
The benchmark results for JumpHash (Algorithm 2) and JumpBackHash* (Algorithm 6) are shown in Figure 3. As reference, we also included non-consistent approaches based on the modulo-operation (1) and a pseudorandom assignment using Lemire’s algorithm (Lemire, 2019). Furthermore, we added a dummy test that just consumed the random keys to make sure that the input is not the bottleneck. To allow a fairer comparison, we used SplitMix (Steele et al., 2014) for generating random values for all approaches that require a PRG.
The results agree with the theoretical runtime analysis and show that JumpBackHash is always faster than JumpHash. Interestingly, the runtime of JumpBackHash is comparable or even faster than the modulo-based approach. The reason is that the modulo operation is a quite expensive operation as it involves an integer division. Therefore, if only uniformity but no monotonicity is needed, it would be much better to use the pseudorandom approach.
4. Future Work
In our tests we used the SplitMix random generator (Steele et al., 2014) which is already quite fast and known to have good quality (Lemire, [n. d.]). However, there might be even faster random generators such as linear congruential random generators (L’Ecuyer, 1999) that have sufficient statistical quality for JumpBackHash.
Another idea, which also needs further investigation, is to use the hashed key directly as the first random value. In this way, one random value computation can be saved and the random generator only needs to be initialized with the key as seed when entering the inner loop. First experiments showed that our statistical tests were still successful when using the SplitMix generator. Of course, this can only work for PRGs where the random values are sufficiently independent of the seed.
5. Conclusion
We have introduced JumpBackHash, a new consistent hash algorithm for distributing keys to buckets. It can be implemented with a few lines of code and requires only a standard random generator as found in standard libraries. As theoretically shown and experimentally confirmed, JumpBackHash has an expected constant runtime and is significantly faster than JumpHash. Its speed also allows safely replacing modulo-based assignments to improve system stability during changes by reducing the probability of spikes in resource consumption.